OpenStudy (anguyennn):

how do you find out what x is: 1/2 x -x/4 >10/1

OpenStudy (anguyennn):

\[\frac{ 1 }{ 2 }x-\frac{ x }{ 4 }>\frac{ 10 }{ 1 }\]

OpenStudy (anonymous):

can one of you help me with my problem please

OpenStudy (welshfella):

first eliminate the fractions by muliplying through by 4 can you do that?

OpenStudy (anguyennn):

why by 4?

OpenStudy (welshfella):

because that will eliminate the fractions 4 is the LCD of 2 and 4.

OpenStudy (anguyennn):

so then would it be 1/8x -2x/8?

OpenStudy (anguyennn):

can someone show me? please

OpenStudy (welshfella):

how did you get that?

OpenStudy (welshfella):